An expandable cube, initially 23 cm on each side contains 2.3 g of helium at 20∘C. 1400 J of heat energy are transferred to this gas.

What is the final volume if the process is at constant pressure? In cm^3
Answer #1

V1 = ( 23 cm )^3 = 12167 cm^3

For Constant Pressure Heat Addition:

---------------------------------------...

T2 = T1 + ( Q ) / ( m ) ( Cp )

T2 = 20.0 C + ( 1400 J ) / ( 2.3 g ) ( 5.1926 J/g - C )

T2 = 137.22 C = 410.22 K

V2 = ( V1 ) ( T2 / T1 )

V2 = ( 12167 cm^3 ) ( 410.22 K / 293 K ) = 17,035 cm^3